# DiRW: Path-Aware Digraph Learning for Heterophily

### v1

@@ 3.1.1. Direction-aware Path Sampler
In DiRW, the direction of edges plays a pivotal role in the sampling process, dictating the trajectory of the random walk. This critical stage is tasked with determining the destination set for each step of the walk, which is essential for calculating the node probabilities and ensuring the continuity of the walk without interruptions.
Specifically, when the walk encounters a node that is devoid of incoming edges, the model strategically steers the walk towards its out-neighborhood. Conversely, if a node lacks outgoing edges, the walk remains confined to its in-neighborhood. This approach effectively circumvents the issue of walk termination at nodes with unidirectional connectivity, which is mentioned in Key Insight 1. For nodes that possess both incoming and outgoing edges, a more nuanced strategy is essential to balance the presence and directionality of edges. Drawing inspiration from the magnetic Laplacian matrix utilized for digraphs (Zhang et al., 2021a), DiRW introduces a hyperparameter . At each walk step, a random value is sampled uniformly from the . If surpasses , the walk proceeds along the directed edge, anchoring the target node within the out-neighbors. Otherwise, the walk considers both the in-neighbors and out-neighbors of the node. The formulation of this process is encapsulated in the following equations:
where is the indicate function, which equals if the condition is met and otherwise, and is the vector of ones. is a scalar that determines the presence of outgoing edges of node . represents the potential destinations matrix guiding the walk towards the next node in the sequence.
3.1.2. Multi-scale and Multi-order Walk Probability
The computation of walk probabilities in DiRW is a nuanced process that seamlessly blends topological structure with node profiles, which allows DiRW to delve into the complexities of both high-order and one-order homophily within heterophilous graphs.
From the topological perspective, DiRW is designed to unearth high-order homophily information which is often embedded within heterophilous graphs (Xie et al., 2023). The model strategically prioritizes nodes which share a limited amount of common neighbors with the current node, bridging to distant regions of the digraph. The topological-based probability measure, denoted as , is crafted to be inversely proportional to the number of common neighbors and normalized by the degree of the candidate node .
Additionally, from the feature standpoint, DiRW captures one-order homophily by favoring walks towards neighbors with more similar features, which present stronger homophily. The feature-based probability is calculated based on the cosine similarity between the features of the current node and candidate node.
The transition probability from node to node in a walk step is determined by aggregating the topological- and feature-based probabilities, which are then normalized by a softmax function.
Recognizing that the softmax scales vary with the walk directions, we pre-compute the topological- and feature-based probabilities ( and ) to obviate repetitive calculations and ensure the operational efficiency of our DiRW.
@@ 3.1.3. Homophily Entropy-based Personalized Walk Length
In DiRW, the walk length is not predefined but is instead dynamically tailored based on according to the homophily entropy. Traditional homophily metrics predominantly focus on immediate neighborhood information, providing a limited view of the walk sequence’s quality. To address this limitation, we introduce a novel metric - homophily entropy - which evaluates the walk sequence quality through the lens of feature similarity among the nodes within the sequence. Our objective is to ensure that each node in the walk sequence maintains a strong homophilous relationship with the originating node, thereby enriching the aggregated information with relevance to it. To achieve this, we first convert the sampled path into a homophily label sequence, represented as follows:
where and denotes the -th element in the walk sequence and the homophily label sequence, respectively. By applying Shannon entropy (Li and Pan, 2016) to it, we derive the homophily entropy:
where represents the probability of observing the -th element in . Our analysis indicates that a lower homophily entropy corresponds to a higher quality walk sequence.
This metric is instrumental in guiding the determination of walk length for each sequence. We initiate by setting a minimum walk length to guarantee the collection of adequate information. Upon reaching , the walk is extended until two consecutive increases in homophily entropy are detected, signifying an encounter with a heterophilous neighborhood. It aligns with the Key Insight 3 outlined in Sec. 1, as homophilous nodes prefer longer walks.
@@ 3.1.4. Weight-free Path Encoding
To account for the varying influence of nodes along a path, DiRW employs an exponential decay function to assign weights and compute the path embedding. This scheme preserves the order information in the sequence and ensures that nodes in proximity to the originating node in the path are accorded greater significance in the path embedding. The weight assigned to each node in the path is calculated as follows:
where is the decay hyperparameter, and denotes the length of the path. The path embedding is obtained by a weighted sum of the node features along the path, given by:
where and represent the -th sampled path and -th path embedding of node , respectively. This weight-free aggregation mechanism allows for the integration of the sampling into the pre-processing stage, which substantially reduces the computational complexity during training. Moreover, the exponential decay mechanism effectively retains the sequential information inherent in the walk sequence, surpassing the constraints of conventional message-passing approaches that often rely on uniform weighting.
@@ 3.1.5. Adaptive Walk Numbers
Ensuring the collection of comprehensive contextual information is critical in DiRW, guiding the determination of the requisite number of walks per node. This is achieved by evaluating the richness of information encapsulated within the path embeddings. The richness is quantified through the average embedding of the sampled paths, represented as:
where is the number of sampled paths. To avoid the premature termination of the walk process, an initial minimum walk number is established. Subsequent to each walk, DiRW calculates the L2 norm of the discrepancy between the current measure of information richness and that of the preceding iteration:
If the value of fall below the predefined threshold , it indicates that the sampling has reached a saturation point, beyond which additional walks are unlikely to contribute substantially to the information richness. In such cases, the walk is discontinued to prevent further unnecessary and redundant sampling, which enhances the efficiency of the information gathering process and is reflective of our Key Insight 2. It acknowledges that heterophilous nodes are often characterized by more intricate neighborhoods and necessitate a greater walk number to amass a sufficiently rich information.
@@ 3.2.1. Attention-based Node Encoding
Following the pre-processing sampling phase, the DiRW model harnesses an attention mechanism to discern and weigh the informativeness of various path embeddings. This process is pivotal for constructing node embeddings that are rich in relevant contextual information.
Specifically, for each node , DiRW initiates the encoding of the sampled paths through a pair of linear layers, interspersed with a LeakyReLU activation function. The resulting encoding is subsequently processed through a softmax function to yield attention scores . The final node embedding is then computed as a weighted aggregation of the path embeddings , with the weights being the attention scores , articulated as follows:
@@ 3.2.2. Node Classifier
With the node embeddings at hand, DiRW deploys an MLP to perform the node classification task, employing the cross-entropy loss function to guide the training process:
where represents the set of labeled training nodes. and correspond to the one-hot encoded true label and the predicted label for node , respectively.
@@ 4. Experiments
In this section, we conduct thorough evaluations of our DiRW. Our experiments are structured to answer four pivotal questions: Q1: Can DiRW achieve superior performance as a new DiGNN and as a plug-and-play strategy for existing PathGNNs? Q2: If DiRW proves to be effective, what factors contribute to its enhanced performance? Q3: What is the running efficiency of DiRW? Q4: How is the path quality of DiRW and other PathGNNs? To maximize the usage for the space, we introduce datasets, baselines, experimental settings in Appendix A.2-A.5, demonstrate the results of link-level tasks in Appendix A.6, and include a sensitivity analysis in Appendix A.7.
@@ 4.1.1. An Innovative Learning Architecture
To answer Q1, we conduct comparative experiments to evaluate the performance of DiRW in node classification. The results presented in Tab. 1 demonstrate that as a novel DiGNN, DiRW achieves exceptional performance across all datasets and outperforms the current leading DiGNN MagNet by an average of 3.2%, which is attributed to its adept modeling of digraphs and heterophilous relationships. In contrast, traditional undirected GNNs perform well in homophilous scenarios but struggle in heterophilous digraphs due to simplistic undirected adjacency matrix approaches. Moreover, while Dir-GNN and ADPA strive to tackle heterophily with digraph modeling, their methods show limitations in effectively generalizing to homophilous digraphs. Furthermore, while undirected PathGNNs show promise in modeling heterophilous graphs, their performance is significantly hindered by critical shortcomings. Notably, these models fail to account for the directionality of edges in digraphs, and their simplistic walk strategies limit their effectiveness in complex digraph structures.
It is imperative to highlight that the sophisticated model architectures such as RNNs employed in PathGNNs engender scalability challenges, leading to out-of-memory (OOM) errors when dealing with large-scale digraphs ogbn-arxiv and ogbn-products. Similarly, the intricate message-passing paradigms that incorporate multiple convolutional layers in DGCN and NSTE result in incomplete training within 12 hours, culminating in out-of-time (OOT) errors. In contrast, DiRW revels in its high efficiency and achieves superior performance on large-scale digraphs with its weight-free path sampling strategy and lightweight learning mechanism.
@@ 4.1.2. A Plug-and-Play Approach
Beyond assessing DiRW as an innovative neural architecture, we also integrate the optimized path sampler with other PathGNNs. Specifically, we combine DiRW with RAWGNN and PathNet to enable modeling of digraphs. Due to the complicated path and node embedding learning rules designed in these two models, it is tough for the weight-free adaptive walking length and number in DiRW to generalize to them. Therefore, we have incorporated only the direction-aware sampling strategy and the walking probabilities that consider both topological structure and node profiles into these two models to explore their performances. The results are demonstrated in Tab. 2.

### v3

@@ 3.1.1. Direction-aware Path Sampler
In DiRW, edge directionality governs the walk trajectory through the direction-aware sampling protocol that dynamically adapts to node connectivity patterns. This critical stage is tasked with determining the destination set for each walk, which is essential for calculating the node probabilities and ensuring the continuity of the walk without interruptions.
Specifically, when the walk encounters a sink node (node with empty in-neighborhoods), it strategically transitions to its out-neighbors. Conversely, when encountering a source node (node lacking out-edges), the walk remains confined to its in-neighborhood. This approach effectively circumvents the issue of walk termination at nodes with unidirectional connectivity, which is mentioned in Key Insight 1. For nodes that possess both incoming and outgoing edges, a more nuanced strategy is needed to balance the edge presence and directionality. Drawing inspiration from the magnetic Laplacian (Zhang et al., 2021a), DiRW introduces a direction control parameter with a uniformly sampled threshold to determine the successor selection. If , the walk proceeds directionally to out-neighbors, whereas permits bidirectional exploration considering both the in-neighbors and out-neighbors. This mechanism interpolates between strict directionality preservation () and undirected-like exploration (), while preventing walk termination at unidirectionally connected nodes.
3.1.2. Multi-order Walk Probability
The computation of walk probabilities in DiRW is a nuanced process that seamlessly integrates both the topological structure and node profiles, enabling DiRW to delve into the complexities of both high-order and one-order homophily within heterophilous graphs.
Topology-based High-order Probability. DiRW uncovers high-order homophily information which is often embedded within heterophilous graphs (Xie et al., 2023) by strategically prioritizing nodes with few common neighbors, thus bridging to distant areas of the graph. The topology-based high-order probability is inversely proportional to the number of common neighbors and normalized by the degree of the candidate node :
Feature-based One-order Probability. From the feature standpoint, DiRW captures one-order homophily by favoring walks towards neighbors with more similar features, which indicate stronger homophily. The feature-based one-order probability is calculated based on the cosine similarity between the features of the current node and candidate node :
The transition probability from node to node in a walk step is determined by combining the topology-based and feature-based probabilities, normalized by a softmax function:
Recognizing that the softmax scales vary with the walk directions, and we pre-compute and to ensure efficiency.
@@ 3.1.3. Homophily Entropy-based Personalized Walk Length
In DiRW, the walk length is dynamically determined based on homophily entropy rather than being predefined and fixed. Traditional homophily metrics predominantly focus on immediate neighborhood information, providing a limited view of the walk sequence’s quality. To address this limitation, we introduce a novel metric Homophily Entropy, which assesses the walk sequence quality based on feature similarity among nodes in the sequence. Our objective is to ensure sampled nodes in the walk sequence have a strong homophilous relationship with the originating node, enriching the aggregated information with relevance to it. We first convert the walk sequence into a homophily label sequence, represented as:
where equals if the condition is met and otherwise. denotes the -th node in the walk sequence, and is the corresponding homophily label. By applying the Shannon entropy (Li and Pan, 2016) on , we derive the homophily entropy:
where represents the probability of encountering in and denotes the length of the sequence. Our analysis indicates that a lower homophily entropy corresponds to a higher walk sequence quality.
This entropy metric serves for dynamically determining node-specific walk lengths. We start with a predefined minimum walk length to guarantee the adequate information acquisition. Subsequent walk persists until two consecutive homophily entropy increments, signaling the heterophilous neighborhood. This adaptive scheme aligns with Key Insight 2, which explicitly accommodates nodes with homophilous neighborhoods.
| Actor | 7,600 | 26,659 | 932 | 5 | 48%/32%/20% | actor network | |
| Rating | 24,492 | 93,050 | 300 | 5 | 50%/25%/25% | rating network |
@@ 3.1.4. Weight-free Path Encoding
To address the varying influence of nodes along a path, DiRW uses an exponential decay function to assign weights and compute the path embedding:
where represents the decay parameter and denotes the -th sampled path of node . This weight-free aggregation mechanism allows for the integration of the sampling into the pre-processing stage, which substantially reduces the computational complexity during training. Moreover, by giving greater significance to nodes closer to the originating node, the exponential decay mechanism effectively captures the sequential information inherent in the walk sequence, overcoming the constraints of uniform weighting often found in message-passing approaches (Kipf and Welling, 2017).
@@ 3.1.5. Adaptive Walk Numbers
In DiRW, gathering comprehensive context is essential, which influences the walk numbers per node. This is determined by assessing the Information Richness, measured by the average of sampled path embeddings. To avoid early walk termination, a minimum walk number is set. After each walk, DiRW computes the L2 norm of the Information Richness difference between the current and previous sampled sequences:
where is the number of sampled path sequences.
If falls below the predefined threshold , it indicates that the sampling has reached a saturation point, beyond which additional walks are unlikely to contribute substantially to the information richness. In such cases, the walk is stopped to avoid further unnecessary and redundant sampling, improving the information gathering efficiency and reflecting Key Insight 2: heterophilous nodes are often characterized by more intricate neighborhoods and necessitate more walks to amass sufficiently rich information.
@@ 3.2.1. Attention-based Node Encoding
After the pre-processing sampling phase, DiRW harnesses an attention mechanism to discern and weigh the informativeness of various path embeddings. This process is pivotal for constructing node embeddings that are rich in relevant contextual information.
Specifically, for each node , DiRW initiates the encoding of the sampled paths through a pair of linear layers, interspersed with a LeakyReLU activation function. The encoded representation is subsequently processed through a softmax function to yield attention scores . The final node embedding is then computed as a weighted aggregation of the path embeddings , with the weights being the attention scores , articulated as follows:
@@ 3.2.2. Node Classifier
With the node embeddings at hand, DiRW deploys an MLP to tackle the node classification task. The training process is directed by the cross-entropy loss as follows:
where represents the training set. and are the one-hot encoded true label and the predicted label of node , respectively.
In the link prediction task, DiRW estimates the probability of edges between node pairs using another MLP. The input to this MLP consists of the concatenated embeddings of the node pairs, allowing the model to leverage the feature information from both nodes. The training process is also guided by the binary cross-entropy loss.
@@ 4. Experiments
In this section, we conduct a comprehensive evaluation of our DiRW, structured to address five critical research dimensions: Q1: Can DiRW establish superior performance as both a new DiGNN and a plug-and-play strategy? Q2: If DiRW proves effective, what architectural characteristics contribute to its enhanced performance? Q3: How does DiRW’s time complexity compare with baseline methods in practical deployment scenarios? Q4: How does the hyperparameters sensitivity and robustness of DiRW manifest? Q5: How does a comparative analysis of path quality between DiRW and existing PathGNNs inform its structural learning advantages?
4.1. Experimental Setup
4.1.1. Datasets.
Our comprehensive evaluation spans 9 benchmark datasets encompassing both directed and undirected graphs with diverse homophily characteristics. We have documented the detailed information of these datasets in the Tab. 1.
4.1.2. Baselines.
Our experiment leverages diverse GNNs as comparative benchmarks, which can be categorized as follows: (i) traditional undirected approaches: GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018); (ii) directed spatial methods: DGCN (Tong et al., 2020b), NSTE (Kollias et al., 2022), DIMPA (He et al., 2022), Dir-GNN (Rossi et al., 2023), ADPA (Sun et al., 2024); (iii) directed spectral methods: DiGCN (Tong et al., 2020a) and its variant DiGCNappr, MagNet (Zhang et al., 2021a), MGC (Zhang et al., 2021b), LightDiC (Li et al., 2024a); (iiii) PathGNNs: RAWGNN (Jin et al., 2022) and PathNet (Sun et al., 2022). To minimize randomness and ensure fair comparisons, we repeated each experiment 10 times to obtain unbiased performances. Moreover, we transform the digraph to undirected graph and feed it into undirected GNNs.
4.1.3. Hyperparameter Settings.
4.1.4. Experimental Environment.
Our experiments are conducted on the machine with Intel(R) Xeon(R) Platinum 8468V, NVIDIA H800 PCIe, and CUDA 12.2. The operating system is Ubuntu 20.04.6. As for the software, we use Python 3.8 and Pytorch 2.2.1.
@@ 4.2.1. An Innovative Learning Architecture
To answer Q1, we conduct comparative experiments to evaluate the performance of DiRW in node classification. The results presented in Tab. 2 demonstrate that, as a novel DiGNN, DiRW achieves exceptional performance across all datasets, outperforming the leading DiGNN, MagNet, by an average of 3.2%. This improvement is attributed to its effective modeling of digraphs and heterophilous relationships.
In contrast, traditional undirected GNNs perform well in homophilous scenarios but struggle in heterophilous digraphs due to their reliance on simplistic undirected adjacency matrix approaches. While Dir-GNN and ADPA attempt to address heterophily through digraph modeling, their methods exhibit limitations in generalizing effectively to homophilous digraphs. Furthermore, although undirected PathGNNs show potential in modeling heterophilous graphs, their performance is significantly compromised by neglect of edge direction and simplistic walk strategies. Notably, they fail to account for edge directionality in digraphs, and their simplistic walk strategies limit their effectiveness in complex digraph structures.
Emphasize that the sophisticated model architectures like RNNs in PathGNNs cause scalability issues, leading to out-of-memory (OOM) errors with large-scale graphs like ogbn-arxiv and ogbn-products. Similarly, the intricate message-passing paradigm with multiple convolutional layers in DGCN and NSTE often leads to incomplete training within 12 hours, resulting in out-of-time (OOT) errors. In contrast, DiRW demonstrates high efficiency and superior performance on large-scale digraphs through its weight-free path sampling strategy and lightweight learning mechanism.
@@ 4.2.2. A Plug-and-Play Approach
In addition to evaluating DiRW as an innovative neural architecture, we also integrate its optimized path sampler with other PathGNNs, specifically combining DiRW with RAWGNN and PathNet for digraph modeling. However, due to the complicated path and node embedding learning rules designed in these two models, it is tough for the weight-free adaptive walking length and number in DiRW to generalize to them. Therefore, we only incorporated the direction-aware sampling strategy and walking probabilities that consider both topological structure and node profiles into these two models to assess their performance. The results are demonstrated in Tab. 3.